On January 1, 2017, Huber Co. sold 12% bonds with a face value of $2,000,000. The bonds mature in five years, and interest is paid semiannually on June 30 and December 31. The bonds were sold for $2,154,500 to yield 10%. Using the effective-interest method of amortization, interest expense for 2017 is what?

**Please explain process. I answered a very similar question, but can not seem to recreate my results using my method.
Answer #1

Answer : $214,836.25 (or) $214,836

The face value of the bond = 2,000,000

Bond price = $ 2,154,500

Bond coupon rate = 12%

Yield rate = 10%

Cash interest = $2,000,000 * 12% *6/12 = $120,000

Interest Expenses = $ 2,154,500*10%*6/12 = $107,725

   = $ 2,142,225*10%*6/12 = $ 107,111.25

Answers : Interst Expenses for 2017 = $ 107,725 + $107111.25 = $214,836.25 (or) $214,836 (Answer)

Period Cash Interest Interest Expenses Bond premium Amortization Bond Carrying value
January 01 2,154,500
June 30 120,000 107,725 12275 2,142,225
December 31 120,000 107,111.25 12889 2,129,336